MULTIPLICATIVE NUMBER THEORY II. MODERN DEVELOPMENTS Multiplicative Number Theory

II. Modern Developments

HUGH L. MONTGOMERY University of Michigan, Ann Arbor

ROBERT C. VAUGHAN Pennsylvania State University, University Park

Hugh L. Montgomery Department of Mathematics University of Michigan Ann Arbor, Michigan 48109–1109

email: hlm@umich.edu

Robert C. Vaughan Department of Mathematics Pennsylvania State University University Park, PA 16802

email: rvaughan@math.psu.edu